OpenStudy (anonymous):

Find the solution for y ' =(y-x)^2

OpenStudy (anonymous):

does that y have a prime sign? or is it simply just y=(y-x)^2

OpenStudy (anonymous):

Yes, y prime equals (y-x)^2 ->y ' = (y-x)^2

OpenStudy (anonymous):

oh ok, i havent learned differential equations so im afraid i cant help you. im sorry.